90 kV + Is 60 kV 30 kV 3 V

Is find your current
Answer #1

Methool-1 Gok2 M YI it Is Eook Bokaz 312 current I 31 30x103 I 30-IMA -3 Ia = 0.1 X10 13 60 - Is 180 > Is = 180 X 0.1X10 60 3+ MW Go !! Method - 2 -90 3 M 30 30 TS ㅎ Apply KCl to the circuit vi-3 tvi Is Чок Gok 2(0, -3) +30 180k => 28-6 +34 Isok 5V,At Node 3, 3-V 3 f gok 3ok => 3-67 +9 :o gok Vi = 3+9=120 ,, So, Is 5X12-6 180k 60-6 Isok Is = 54 180k Is= 0.3mA